# Fast Policy Learning for Linear Quadratic Regulator with Entropy Regularization

### v1

@@ Abstract
This paper proposes and analyzes two new policy learning methods: regularized policy gradient (RPG) and iterative policy optimization (IPO), for a class of discounted linear-quadratic regulator (LQR) problems over an infinite time horizon with entropy regularization. Assuming access to the exact policy evaluation, both proposed approaches are proved to converge linearly in finding optimal policies of the regularized LQR. Moreover, the IPO method can achieve a super-linear convergence rate once it enters a local region around the optimal policy. Finally, when the optimal policy from a well-understood environment in an RL problem is appropriately transferred as the initial policy to an RL problem with an unknown environment, the IPO method is shown to enable a super-linear convergence rate if the latter is sufficiently close to the former. The performances of these proposed algorithms are supported by numerical examples.
@@ 1 Introduction
This paper proposes and analyzes two new policy learning methods: regularized policy gradient (RPG) and iterative policy optimization (IPO), for a class of discounted entropy-regularized linear-quadratic regulator (LQR) problems over an infinite time horizon. Assuming access to the exact policy evaluation, both proposed approaches are shown to converge linearly in finding optimal policies of the regularized LQR (Theorem 4.1 and 5.1). Moreover, the IPO method can achieve a super-linear convergence rate (on the order of one and a half) once it enters a local region around the optimal policy. Finally, when the optimal policy from a well-understood environment in an RL problem is appropriately transferred as the initial policy to an RL problem with an unknown environment, the IPO algorithm is shown to enable a super-linear convergence rate if the latter is sufficiently close to the former (Theorem 6.1).
Our analysis approach is inspired by [12] to establish the gradient dominance condition within the linear-quadratic structure. Unlike theirs, our framework incorporates entropy regularization and state transition noise (Section 2). Therefore, in contrast to their deterministic and linear policies, our policies are of Gaussian type. Consequently, the gradient dominance condition involves both the gradient of the mean and the gradient of the covariance (Lemma 3.2). Accordingly, to establish the convergence of the covariance update in RPG, the smoothness of the objective function for bounded covariance is exploited, which is ensured with proper initialization (Lemma 4.1).
@@
Related works of policy gradient methods in LQR.
As a cornerstone in optimal control theory, the LQR problem is to find an optimal control in a linear dynamical system with a quadratic cost. LQR is popular due to its analytical tractability and its approximation power to nonlinear problems [3]. Until recently, most works on the LQR problem assumed that the model parameters are fully known. The first global convergence result for the policy gradient method to learn the optimal policy for LQR problems was developed in [12] for an infinite time horizon and with deterministic dynamics. Their work was extended in [6] to give global optimality guarantees of policy gradient methods for a larger class of control problems, which satisfy a closure condition under policy improvement and convexity of policy improvement steps. More progress has been made for policy gradient methods in other settings as well, including [7] for a real-valued matrix function, [8] for a continuous-time setting, [13] for multiplicative noise, [23], [25] for additive noise, and [17] for finite time horizon with an additive noise, [31] and [38] for time-average costs with risk constraints, [19] for nearly-linear dynamic system, and [34] for distributional LQR to find the distribution of the return. Our work establishes (fast) convergence for both policy gradient based and policy optimization based algorithms for an infinite time horizon LQR with entropy regularization.
@@
Entropy regularization has been frequently adopted to encourage exploration and improve convergence [15, 20, 21, 22, 28, 29, 32, 33, 36]. In particular, [2] showed that entropy regularization induces a smoother landscape that allows for the use of larger learning rates, and hence, faster convergence. Convergence rate analysis has been established when the underlying dynamic is an MDP with finite states and finite actions. For instance, [1] and [26] developed convergence guarantees for regularized policy gradient methods, with relative entropy regularization considered in [1] and entropy regularization in [26]. Both papers suggest the role of regularization in guaranteeing faster convergence for the tabular setting. For the natural policy gradient method, [11] established a global linear convergence rate and a local quadratic convergence rate.
For system with infinite number of states and actions, the closest to our work in terms of model setup is [33]. Our work replaces their aggregated control setup with controls that are randomly sampled from the policy, which are more realistic in handling real-world systems. The focuses of these two papers are also different: theirs explained the exploitation–exploration trade-off with entropy regularization from a continuous-time stochastic control perspective and provided theoretical support for Gaussian exploration for LQR; while ours is on algorithms design and the convergence analysis. To the best of our knowledge, our work is the first non-asymptotic convergence result for LQR under entropy regularization.
@@
The rest of the paper is organized as follows. Section 2 introduces the problem set-up and provides its theoretical solution using the dynamic programming principle. Section 3 presents the gradient dominance condition and related smoothness property. Section 4 introduces the RPG method and provides the global linear convergence result, and Section 5 proposes the IPO method, along with its global linear convergence and local super-linear convergence results. Finally, Section 6 shows that IPO leads to an efficient transfer learning scheme. Numerical examples are given in Section 7.
Throughout the paper, unless otherwise specified, proofs of technical lemmas are left in the Appendix.
2 Regularized LQR Problem and Solution
@@ 2.1 Problem Formulation
We consider an entropy-regularized LQR problem over an infinite time horizon with a constant discounted rate.
@@
$J_{\pi}(x):=\mathbb{E}_{\pi}\left.\left[\sum_{t=1}^{\infty}\gamma^{t}\left(x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}+\tau\log\pi(u_{t}|x_{t})\right)\right|x_{0}=x\right].$ (2)
such that for $t=0,1,2,\cdots$,
@@ 2.2 Optimal Value Function and Policy
While the optimal solution to the Linear Quadratic Regulator (LQR) problem is a well-explored topic, it is worth noting that, to the best of our knowledge, no prior work has presented a solution to the entropy-regularized LQR problem in the form of (2). Additionally, in the study of [33] for entropy-regularized LQR with continuous-time state dynamics, they focused on the state transitions with aggregated controls. This differs from the state transitions considered in (3), where the controls in the state transitions are randomly sampled from the policy $\pi$.
@@ 4 Regularized Policy Gradient Method
Remark 4.1 (Comparison to natural policy gradient.).
@@ 5 Iterative Policy Optimization Method
From (17), one can obtain the value function starting at $x$ with parameter $(K,\Sigma):$
$J_{K,\Sigma}(x)=$ $x^{\top}(Q+K^{\top}RK)x+\operatorname{Tr}(\Sigma R)-\frac{\tau}{2}\left(k+\log\left((2\pi)^{k}\operatorname{det}\Sigma\right)\right)$
$+\gamma x^{\top}(A-BK)^{\top}P_{K}(A-BK)x+\gamma\operatorname{Tr}(WP_{K})+\gamma\operatorname{Tr}(\Sigma B^{\top}P_{K}B).$
@@
$K^{\prime}=\arg\max_{\widetilde{K}}\,$ $x^{\top}(Q+\widetilde{K}^{\top}R\widetilde{K})x+\operatorname{Tr}(\Sigma R)-\frac{\tau}{2}\left(k+\log\left((2\pi)^{k}\operatorname{det}\Sigma\right)\right)$
$+\gamma x^{\top}(A-B\widetilde{K})^{\top}P_{K}(A-B\widetilde{K})x+\gamma\operatorname{Tr}(WP_{K})+\gamma\operatorname{Tr}(\Sigma B^{\top}P_{K}B),$
and
$\Sigma^{\prime}=\arg\max_{\widetilde{\Sigma}}\,\operatorname{Tr}(\widetilde{\Sigma}R)-\frac{\tau}{2}\left(k+\log\left((2\pi)^{k}\operatorname{det}\widetilde{\Sigma}\right)\right)+\gamma\operatorname{Tr}(\widetilde{\Sigma}B^{\top}P_{K}B).$
@@ 5.1 Global Linear Convergence
Theorem 5.1 suggests that (IPO) achieves a global linear convergence. Compared with (RPG), (IPO) exhibits faster convergence in terms of the rate at which the objective function decreases (cf. Lemmas 4.2 and 5.1). Furthermore, the subsequent section demonstrates that (IPO) enjoys a super-linear local convergence when the initial policy parameter is within a neighborhood of the optimal policy parameter.
@@ 5.2 Super-linear Local Convergence
$\leq$ $\left(\xi_{\gamma,\rho}\|\mathbb{E}[x_{0}x_{0}^{\top}]\|+\zeta_{\gamma,\rho}\|B\Sigma B^{\top}+W\|\right)\cdot 2\rho\|B\|\left\|K_{1}-K_{2}\right\|+\omega_{\gamma,\rho}\|B\|^{2}\|\Sigma_{1}-\Sigma_{2}\|.$
@@ 6 Application: Transfer Learning
We analyze two environments $\mathcal{M}:=(A,B)$ and ¯$\overline{\mathcal{M}}:=(\overline{A},\overline{B})$, with $(K^{*},\Sigma^{*})$ and $(\overline{K}^{*},\overline{\Sigma}^{*})$ their respective optimal policies. Assume that one has access to the optimal (regularized) policy $(K^{*},\Sigma^{*})$ for environment $\mathcal{M}$, called the well-understood environment. We use $(K^{*},\Sigma^{*})$ as a policy initialization for the less understood environment $\overline{\mathcal{M}}$, called the new environment. The goal is to investigate under what conditions this initialization enters the super-linear convergence regime of $\overline{\mathcal{M}}$.
@@ 7 Numerical Experiments
Appendix A Proofs in Section 2
@@
where $\mathcal{L}(u,p,\lambda):=(u^{\top}Mu+b^{\top}u)p+\tau p\log p+\lambda p.$ Since $\frac{\partial\mathcal{L}}{\partial{p}}=\lambda+u^{\top}Mu+b^{\top}u+\tau+\tau\log p,$ by setting $\frac{\partial\mathcal{L}}{\partial{p}}=0,$ we then get $p(u)\propto e^{-\frac{1}{\tau}(u^{\top}Mu+b^{\top}u)}.$ Thus, the optimal solution $p^{*}$ is a multivariate Gaussian distribution with $\mathcal{N}\left(-\frac{1}{2}M^{-1}b,\frac{\tau}{2}M^{-1}\right)$.
Appendix B Proofs in Section 3
B.1 Proof of Lemma 3.1
$\nabla_{\Sigma}C(K,\Sigma)$ in (18) can be checked by direct gradient calculation. To verify (18), first fix $K\in\mathbb{R}^{k\times n}$ and define $f:\mathcal{X}\rightarrow\mathbb{R}$ by $f(y):=y^{\top}P_{K}y,\forall y\in\mathcal{X}$ and we aim to find the gradient of $f$ with respect to $K$. For any $y\in\mathcal{X},$ by the Riccati Equation for $P_{K}$ in (17) we have
$f(y)$ $=y^{\top}\left(\gamma(A-BK)^{\top}P_{K}(A-BK)+Q+K^{\top}RK\right)y$
$=\gamma f\left((A-BK)y\right)+y^{\top}\left(Q+K^{\top}RK\right)y.$
$\nabla_{K}f((A-BK)y)$ has two terms: one with respect to the input $(A-BK)y$ and one with respect to $K$ in the subscript of $P_{K}$. This implies
$\nabla_{K}f(y)$ $=2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)yy^{\top}+\gamma\nabla_{K}f(y^{\prime})|_{y^{\prime}=(A-BK)y}$ (41)
@@ Appendix B Proofs in Section 3
$\nabla_{K}C(K,\Sigma)=$ $\mathbb{E}[\nabla_{K}f(x_{0})]+\nabla_{K}q_{K,\Sigma}$
=$\overset{(a)}{=}$ $\mathbb{E}\Big{[}2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)x_{0}x_{0}^{\top}+\gamma\nabla_{K}f(\hat{x}_{1})|_{\hat{x}_{1}=(A-BK)x_{0}}$
@@
=$\overset{(b)}{=}$ $\mathbb{E}\Big{[}2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)x_{0}x_{0}^{\top}+\gamma\nabla_{K}f({x}_{1})|_{{x}_{1}=A+Bu_{0}+w_{0}}$
@@
where $(a)$ follows from applying (41) with $y=x_{0}$ and taking the gradient of $q_{K}$ in (17) with respect to $K$; $(b)$ follows from
$\mathbb{E}_{x_{0},u_{0},w_{0}}[f(Ax_{0}+Bu_{0}+w_{0})]=\mathbb{E}_{x_{0},w_{0}}[f\left((A-BK)x_{0}\right)+w_{0}^{\top}P_{K}w_{0}]+\Sigma B^{\top}P_{K}B.$
@@
With the notations of the Q-function in (42) and the advantage function in (44), we first provide a convenient form for the difference of the cost functions with respect to two different policies in Lemma B.1. This will be used in the proof of Lemma 3.2.
Lemma B.1 (Cost difference).
Suppose policies $\pi$ and $\pi^{\prime}$ are in form of (14) with parameters $(K^{\prime},\Sigma^{\prime})$ and $(K,\Sigma).$ Let $\{x_{t}^{\prime}\}$ and $\{u_{t}^{\prime}\}$ be state and action sequences generated by $\pi^{\prime}$ with noise sequence $\{w_{t}^{\prime}\},$ i.e., $x_{t+1}^{\prime}=Ax_{t}^{\prime}+Bu_{t}^{\prime}+w_{t}^{\prime}$. Then
$J_{K^{\prime},\Sigma^{\prime}}(x)-J_{K,\Sigma}(x)=\mathbb{E}_{\pi^{\prime}}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{K,\Sigma}(x_{t}^{\prime},u_{t}^{\prime})\right],$ (45)
where the expectation is taken over $u_{t}^{\prime}\sim\pi^{\prime}(\cdot|x_{t}^{\prime})$ for all $t\geq 0$ and i.i.d noise $\{w_{t}^{\prime}\}_{t=0}^{\infty}$ with mean 00 and covariance $W$.
@@
(of Lemma B.1). Let $\{c_{t}^{\prime}\}$ be the cost sequences generated by $\pi^{\prime},$ i.e., $c_{t}^{\prime}=x_{t}^{\prime\top}Qx_{t}^{\prime}+u_{t}^{\prime\top}Ru_{t}^{\prime}+\tau\log\pi^{\prime}(u_{t}^{\prime}|x_{t}^{\prime}),$ and $J_{K^{\prime},\Sigma^{\prime}}(x)=\mathbb{E}_{\pi^{\prime}}\left[\sum_{t=0}^{\infty}\gamma^{t}c_{t}^{\prime}\right].$ Then we have
@@
where $(a)$ follows from $x_{0}^{\prime}=x$By (42) and (44), we can continue the above calculations by
@@
(of Lemma 3.2). By Lemma B.1, for any $\pi$ and $\pi^{\prime}$ in form of (14) with parameter $(K,\Sigma)$ and $(K^{\prime},\Sigma^{\prime})$ respectively, then for any $x\in\mathcal{X}$:
$\mathbb{E}_{\pi^{\prime}}\left[A_{K,\Sigma}(x,u)\right]=(1-\gamma)(f_{K}(\Sigma)-f_{K}(\Sigma^{\prime}))-\operatorname{Tr}\left(xx^{\top}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)$
@@
with equality when $K-(R+\gamma B^{\top}P_{K}B)^{-1}E_{K}=K^{\prime}$ holds.
Let $\{x_{t}^{*}\}_{t=0}^{\infty}$ and $\{u_{t}^{*}\}_{t=0}^{\infty}$ be the state and control sequences generated under $\pi^{*}(\cdot|x)=\mathcal{N}(-K^{*}x,\Sigma^{*})$ with noise sequence $\{w_{t}^{*}\}_{t=0}^{\infty}$ with mean 00 and covariance $W$. Apply Lemma B.1 and (47) with $\pi$ and $\pi^{*}$ to get:
$C(K,\Sigma)-C(K^{*},\Sigma^{*})=-\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{K,\Sigma}(x_{t}^{*},u_{t}^{*})\right]$
$\leq\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)+\left(f_{K}(\Sigma^{*})-f_{K}(\Sigma)\right),$ (48)
@@
$\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)\leq\frac{\|S_{K^{*},\Sigma^{*}}\|}{\sigma_{\min}(R)}\operatorname{Tr}(E_{K}^{\top}E_{K}).$ (49)
We aim to bound $\operatorname{Tr}(E_{K}^{\top}E_{K})$ with $\|\nabla_{K}C(K,\Sigma)\|_{F}.$ Apply (18) to get
$\operatorname{Tr}(\nabla_{K}^{\top}C(K,\Sigma)\nabla_{K}C(K,\Sigma))$ $=4\operatorname{Tr}\left(S_{K,\Sigma}E_{K}^{\top}E_{K}S_{K,\Sigma}\right)\geq 4\mu^{2}\operatorname{Tr}(E_{K}^{\top}E_{K}).$
Plug the above result into (49) and use (50) to get (19).
To analyze the second terms in (48), note that $f_{K}$ is a concave function, thus we can find its maximizer $\Sigma_{K}^{*}$ by taking the gradient of $f_{K}$ and setting it to 00 such that $\Sigma_{K}^{*}=\frac{\tau}{2}(R+\gamma B^{\top}P_{K}B)^{-1}.$ Thus,

### v3

# Fast Policy Learning for Linear-Quadratic Control with Entropy Regularization

@@ Abstract
This paper proposes and analyzes two new policy learning methods: regularized policy gradient (RPG) and iterative policy optimization (IPO), for a class of discounted linear-quadratic control (LQC) problems over an infinite time horizon with entropy regularization. Assuming access to the exact policy evaluation, both proposed approaches are proved to converge linearly in finding optimal policies of the regularized LQC. Moreover, the IPO method can achieve a super-linear convergence rate once it enters a local region around the optimal policy. Finally, when the optimal policy for an RL problem with a known environment is appropriately transferred as the initial policy to an RL problem with an unknown environment, the IPO method is shown to enable a super-linear convergence rate if the two environments are sufficiently close. Performances of these proposed algorithms are supported by numerical examples.
@@ 1 Introduction
This paper proposes and analyzes two new policy learning methods: regularized policy gradient (RPG) and iterative policy optimization (IPO), for a class of discounted entropy-regularized linear-quadratic control (LQC) problems over an infinite time horizon. Assuming access to the exact policy evaluation, both approaches are shown to converge linearly in finding optimal policies of the regularized LQC (Theorem 4.2 and 5.1). Moreover, the IPO method can achieve a super-linear convergence rate (on the order of one and a half) once it enters a local region around the optimal policy. Finally, when the optimal policy from an RL problem with a known environment is appropriately transferred as the initial policy to an RL problem with an unknown environment, the IPO algorithm is shown to enable a super-linear convergence rate if the two environments are sufficiently close (Theorem 6.1).
Our analysis approach is inspired by [12] to establish the gradient dominance condition within the linear-quadratic structure. Unlike theirs, our framework incorporates entropy regularization and state transition noise (Section 2). Therefore, in contrast to their deterministic and linear policies, our policies are of Gaussian type. Consequently, the gradient dominance condition involves both the gradient of the mean and the gradient of the covariance (Lemma 3.2). Accordingly, to establish the convergence of the covariance update in RPG, the smoothness of the objective function for bounded covariance is exploited, which is ensured with proper learning rate (Lemma 4.3).
@@
Related works of policy gradient methods in LQC.
As a cornerstone in optimal control theory, the LQC problem is to find an optimal control in a linear dynamical system with a quadratic cost. LQC is popular due to its analytical tractability and its approximation power to nonlinear problems [3]. Until recently, most works on the LQC problem assumed that the model parameters are fully known. The first global convergence result for the policy gradient method to learn the optimal policy for LQC problems was developed in [12] for an infinite time horizon and with deterministic dynamics. Their work was extended in [6] to give global optimality guarantees of policy gradient methods for a larger class of control problems, which satisfy a closure condition under policy improvement and convexity of policy improvement steps. More progress has been made for policy gradient methods in other settings as well, including [7] for a real-valued matrix function, [8] for a continuous-time setting, [14] for multiplicative noise, [25], [27] for additive noise, and [18] for finite time horizon with an additive noise, [37] and [44] for time-average costs with risk constraints, [21] for nearly-linear dynamic system, [40] for distributional LQC to find the distribution of the return, and [20] for nonlinear stochastic control with exit time. Our work establishes fast convergence for both policy gradient based and policy optimization based algorithms for an infinite time horizon LQC with entropy regularization.
@@
Entropy regularization has been frequently adopted to encourage exploration and improve convergence [16, 22, 23, 24, 31, 32, 38, 39, 42, 36, 35, 34, 28]. In particular, [2] showed that entropy regularization induces a smoother landscape that allows for the use of larger learning rates, and hence, faster convergence. Convergence rate analysis has been established when the underlying dynamic is an MDP with finite states and finite actions. For instance, [1] and [29] developed convergence guarantees for regularized policy gradient methods, with relative entropy regularization considered in [1] and entropy regularization in [29]. Both papers suggest the role of regularization in guaranteeing faster convergence for the tabular setting. For the natural policy gradient method, [11] established a global linear convergence rate and a local quadratic convergence rate.
For system with infinite number of states and actions, the closest to our work in terms of model setup is [39]. Our work replaces their aggregated control setup with controls that are randomly sampled from the policy, which are more realistic in handling real-world systems. The focuses of these two papers are also different: theirs explained the exploitation–exploration trade-off with entropy regularization from a continuous-time stochastic control perspective and provided theoretical support for Gaussian exploration for LQC; while ours is on algorithms design and the convergence analysis for discrete-time dynamics. It is worth noting that recently [13] showed the linear convergence rate of the policy gradient method under the framework of [39] with aggregated controls in continuous time. To the best of our knowledge, our work is the first non-asymptotic convergence result for discrete-time LQC with randomized controls under entropy regularization. In addition, we establish the first local super-linear convergence result under the LQC setting.
@@
The rest of the paper is organized as follows. Section 2 introduces the problem and provides its theoretical solution using the dynamic programming principle. Section 3 presents the gradient dominance condition and related smoothness property. Section 4 introduces the RPG method and provides the global linear convergence result, and Section 5 proposes the IPO method, along with its global linear convergence and local super-linear convergence results. Section 6 shows that IPO leads to an efficient transfer learning scheme. Numerical examples are presented in Section 7.
Throughout this paper, proofs of main technical lemmas, unless otherwise specified, are deferred to Section 8.
2 Regularized LQC Problem and Solution
@@ 2.1 Problem Formulation
We consider an entropy-regularized LQC problem over an infinite time horizon with a constant discounted rate.
@@
$J_{\pi}(x):=\mathbb{E}_{\pi}\left.\left[\sum_{t=0}^{\infty}\gamma^{t}\left(x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}+\tau\log\pi(u_{t}|x_{t})\right)\right|x_{0}=x\right],$ (2)
and for $t=0,1,2,\cdots$,
@@ 2.2 Optimal Value Function and Policy
While the optimal solution to the LQC problem is a well-explored topic, it is worth noting that, to the best of our knowledge, no prior work has presented a solution to the entropy-regularized LQC problem in the form of (2). Additionally, in the study of [39] for entropy-regularized LQC with continuous-time state dynamics, they focused on the state transitions with aggregated controls. This differs from the state transitions considered in (3), where the controls in the state transitions are randomly sampled from the policy $\pi$.
@@ 4 Regularized Policy Gradient Method
Remark 4.1 (Comparison to natural policy gradient).
@@ 5 Iterative Policy Optimization Method
By Bellman equation for the value function $J_{K,\Sigma}$,
$J_{K,\Sigma}(x)=\mathbb{E}_{u\sim\pi_{K,\Sigma}}\Big{[}$ $x^{\top}Qx+u^{\top}Ru+\tau\log\pi_{K,\Sigma}(u|x)+\gamma J_{K,\Sigma}(Ax+Bu+w)\Big{]}.$
@@
$K^{\prime},\Sigma^{\prime}=\arg\min_{\widetilde{K},\widetilde{\Sigma}}\mathbb{E}_{u\sim\pi_{\widetilde{K},\widetilde{\Sigma}}}\Big{[}x^{\top}Qx+u^{\top}Ru+\tau\log\pi_{\widetilde{K},\widetilde{\Sigma}}(u|x)+\gamma{J_{K,\Sigma}}(Ax+Bu+w)\Big{]}$
@@ 5.1 Global Linear Convergence
Theorem 5.1 suggests that (IPO) achieves a global linear convergence. Compared with (RPG), (IPO) exhibits faster convergence in terms of the rate at which the objective function decreases (cf. Lemmas 4.4 and 5.2). Furthermore, the subsequent section demonstrates that (IPO) enjoys a super-linear local convergence when the initial policy parameter is within a neighborhood of the optimal policy parameter.
@@ 5.2 Super-linear Local Convergence
$\leq$ $\left(\xi_{\gamma,\rho}\|\mathbb{E}[x_{0}x_{0}^{\top}]\|+\zeta_{\gamma,\rho}\|B\Sigma_{1}B^{\top}+W\|\right)\cdot 2\rho\|B\|\left\|K_{1}-K_{2}\right\|+\omega_{\gamma,\rho}\|B\|^{2}\|\Sigma_{1}-\Sigma_{2}\|.$
@@ 6 Application: Transfer Learning
We analyze two environments $\mathcal{M}:=(A,B)$ and ¯$\overline{\mathcal{M}}:=(\overline{A},\overline{B})$, with $(K^{*},\Sigma^{*})$ and $(\overline{K}^{*},\overline{\Sigma}^{*})$ as their respective optimal policies. Assume that one has access to the optimal (regularized) policy $(K^{*},\Sigma^{*})$ for environment $\mathcal{M}$, called the well-understood environment. We use $(K^{*},\Sigma^{*})$ as a policy initialization for the less understood environment $\overline{\mathcal{M}}$, called the new environment. The goal is to investigate under what conditions this initialization enters the super-linear convergence regime of $\overline{\mathcal{M}}$.
@@ 7 Numerical Experiments
8 Detailed Proofs
8.1 Proofs in Section 2
@@
where $\mathcal{L}(u,v,\lambda):=(u^{\top}Mu+b^{\top}u)v+\tau v\log v+\lambda v.$ Since $\frac{\partial\mathcal{L}}{\partial{v}}=\lambda+u^{\top}Mu+b^{\top}u+\tau+\tau\log v,$ by setting $\frac{\partial\mathcal{L}}{\partial{v}}=0,$ we then get $v\propto e^{-\frac{1}{\tau}(u^{\top}Mu+b^{\top}u)}.$ Thus, the optimal solution $p^{*}$ is a multivariate Gaussian distribution with $\mathcal{N}\left(-\frac{1}{2}M^{-1}b,\frac{\tau}{2}M^{-1}\right)$.
8.2 Proofs in Section 3
8.2.1 Proof of Lemma 3.1
$\nabla_{\Sigma}C(K,\Sigma)$ in (18) can be checked by direct gradient calculation. To verify $\nabla_{K}C(K,\Sigma)$ in (18), first define $f:\mathcal{X}\times\mathbb{R}^{k\times n}\rightarrow\mathbb{R}$ by $f(y,K):=y^{\top}P_{K}y,\forall y\in\mathcal{X}$ and we aim to find the gradient of $f$ with respect to $K$. For any $y\in\mathcal{X},$ by the Riccati Equation for $P_{K}$ in (17) we have
$f(y,K)$ $=y^{\top}\left(\gamma(A-BK)^{\top}P_{K}(A-BK)+Q+K^{\top}RK\right)y$
$=\gamma f\left((A-BK)y,K\right)+y^{\top}\left(Q+K^{\top}RK\right)y.$
$\nabla_{K}f((A-BK)y,K)$ has two terms: one with respect to the input $(A-BK)y$ and one with respect to $K$ in the subscript of $P_{K}$. This implies
$\nabla_{K}f(y,K)=2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)yy^{\top}+\gamma\nabla_{K}f(y^{\prime},K)|_{y^{\prime}=(A-BK)y}$ (41)
@@ 8.2 Proofs in Section 3
$\nabla_{K}C(K,\Sigma)=$ $\mathbb{E}[\nabla_{K}f(x_{0},K)]+\nabla_{K}q_{K,\Sigma}$
=$\overset{(a)}{=}$ $\mathbb{E}\Big{[}2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)x_{0}x_{0}^{\top}+\gamma\nabla_{K}f(\hat{x}_{1},K)|_{\hat{x}_{1}=(A-BK)x_{0}}$
@@
=$\overset{(b)}{=}$ $\mathbb{E}\Big{[}2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)x_{0}x_{0}^{\top}+\gamma\nabla_{K}f({x}_{1},K)|_{{x}_{1}=A+Bu_{0}+w_{0}}$
@@
where $(a)$ follows from applying (41) with $y=x_{0}$ and taking the gradient of $q_{K,\Sigma}$ in (17) with respect to $K$; $(b)$ follows from
$\mathbb{E}_{x_{0},u_{0},w_{0}}[f(Ax_{0}+Bu_{0}+w_{0}),K]=\mathbb{E}_{x_{0},w_{0}}[f\left((A-BK)x_{0},K\right)+w_{0}^{\top}P_{K}w_{0}]+\Sigma B^{\top}P_{K}B.$
@@
With the notations of the Q-function in (42) and the advantage function in (44), we first provide a convenient form for the difference of the cost functions with respect to two different policies in Lemma 8.1. This will be used in the proof of Lemma 3.2.
Lemma 8.1 (Cost difference).
Suppose policies $\pi$ and $\pi^{\prime}$ are in form of (14) with parameters $(K^{\prime},\Sigma^{\prime})$ and $(K,\Sigma).$ Let $\{x_{t}^{\prime}\}_{t=0}^{\infty}$ and $\{u_{t}^{\prime}\}_{t=0}^{\infty}$ be state and action sequences generated by $\pi^{\prime}$ with noise sequence $\{w_{t}^{\prime}\}_{t=0}^{\infty}$ (i.i.d with mean 00 and covariance $W$), i.e., $x_{t+1}^{\prime}=Ax_{t}^{\prime}+Bu_{t}^{\prime}+w_{t}^{\prime}$. Then for any $x\in\mathcal{X},$
$J_{K^{\prime},\Sigma^{\prime}}(x)-J_{K,\Sigma}(x)=\mathbb{E}_{\pi^{\prime}}\left[\left.\sum_{t=0}^{\infty}\gamma^{t}A_{K,\Sigma}(x_{t}^{\prime},u_{t}^{\prime})\right|x_{0}^{\prime}=x\right],$ (45)
where the expectation is taken over $u_{t}^{\prime}\sim\pi^{\prime}(\cdot|x_{t}^{\prime})$ and $w_{t}$ for all $t=0,1,2,\cdots$.
@@
(of Lemma 8.1). Let $\{c_{t}^{\prime}\}_{t=0}^{\infty}$ be the cost sequences generated by $\pi^{\prime},$ i.e., $c_{t}^{\prime}=x_{t}^{\prime\top}Qx_{t}^{\prime}+u_{t}^{\prime\top}Ru_{t}^{\prime}+\tau\log\pi^{\prime}(u_{t}^{\prime}|x_{t}^{\prime}),$ and $J_{K^{\prime},\Sigma^{\prime}}(x)=\mathbb{E}_{\pi^{\prime}}\left[\sum_{t=0}^{\infty}\gamma^{t}c_{t}^{\prime}\right]$. Then we have
@@
where $(a)$ follows from $x_{0}^{\prime}=x$. By (42) and (44), we can continue the above calculations by
@@
(of Lemma 3.2). By Lemma 8.1, for any $\pi$ and $\pi^{\prime}$ in form of (14) with parameter $(K,\Sigma)$ and $(K^{\prime},\Sigma^{\prime})$ respectively, and for any $x\in\mathcal{X}$,
$\mathbb{E}_{\pi^{\prime}}[A_{K,\Sigma}$ $(x,u)]=(1-\gamma)(f_{K}(\Sigma)-f_{K}(\Sigma^{\prime}))-\operatorname{Tr}\left(xx^{\top}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)$
@@
with equality when $K-(R+\gamma B^{\top}P_{K}B)^{-1}E_{K}=K^{\prime}$ holds. Let $\{x_{t}^{*}\}_{t=0}^{\infty}$ and $\{u_{t}^{*}\}_{t=0}^{\infty}$ be the state and control sequences generated under $\pi^{*}(\cdot|x)=\mathcal{N}(-K^{*}x,\Sigma^{*})$ with noise sequence $\{w_{t}^{*}\}_{t=0}^{\infty}$ with mean 00 and covariance $W$. Apply Lemma 8.1 and (47) with $\pi$ and $\pi^{*}$ to get:
$C(K,\Sigma)-C(K^{*},\Sigma^{*})=-\mathbb{E}_{\pi^{*}}\Bigg{[}\sum_{t=0}^{\infty}\gamma^{t}A_{K,\Sigma}(x_{t}^{*},u_{t}^{*})\Bigg{]}$
$\leq\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)+f_{K}(\Sigma^{*})-f_{K}(\Sigma),$ (48)
@@
$\operatorname{Tr}\left(S_{K^{*},\Sigma^{*}}E_{K}^{\top}\left(R+\gamma B^{\top}P_{K}B\right)^{-1}E_{K}\right)\leq\frac{\|S_{K^{*},\Sigma^{*}}\|}{\sigma_{\min}(R)}\operatorname{Tr}(E_{K}^{\top}E_{K})$ (49)
$\leq\frac{\|S_{K^{*},\Sigma^{*}}\|}{\mu^{2}\sigma_{\min}(R)}\operatorname{Tr}\left(S_{K,\Sigma}E_{K}^{\top}E_{K}S_{K,\Sigma}\right)=\frac{\|S_{K^{*},\Sigma^{*}}\|}{4\mu^{2}\sigma_{\min}(R)}\operatorname{Tr}(\nabla_{K}^{\top}C(K,\Sigma)\nabla_{K}C(K,\Sigma)),$
where the last equation follows from (18).
To analyze the second terms in (48), note that $f_{K}$ is a concave function, thus we can find its maximizer $\Sigma_{K}^{*}$ by taking the gradient of $f_{K}$ and setting it to 00, i.e., $\Sigma_{K}^{*}=\frac{\tau}{2}(R+\gamma B^{\top}P_{K}B)^{-1}.$ Thus,